OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

Im thinking hes 13

OpenStudy (anonymous):

10 + 3 is 13!

OpenStudy (anonymous):

Ummm im confused

OpenStudy (anonymous):

X-3=10+X=2(X-3)

OpenStudy (anonymous):

?